a rectangular porch has dimensions of (x+3)feet and (x+4)feet. The area of the porch is 132 square feet. What are the length and width of the porch?
Answer by stanbon(75887)

a rectangular porch has dimensions of (x+3)feet and (x+4)feet. The area of the porch is 132 square feet. What are the length and width of the porch?
---------------------
Equation:
(x+3)(x+4) = 132
------
x^2 + 7x + 12 = 132
-----
x^2 + 7x - 120 = 0
--------
(x-8)(x+15) = 0
-----
x = 8
------
Dimensions:
x+3 = 11
x + 4 = 12
------------
Cheers,
Stan H.
--------